# On second-order variational analysis of variational convexity of prox-regular functions

### v1

@@
Abstract. Variational convexity, together with ist strong counterpart, of extended-real-valued functions has been recently introduced by Rockafellar. In this paper we present second-order characterizations of these properties, i.e., conditions using first-order generalized derivatives of the subgradient mapping. Up to now, such characterizations are only known under the assumptions of prox-regularity and subdifferential continuity and in this paper we discard the latter. To this aim we slightly modify the definitions of the generalized derivatives to be compatible with the $f$-attentive convergence appearing in the definition of subgradients. We formulate our results in terms of both coderivatives and subspace containing derivatives. We also give formulas for the exact bound of variational convexity and study relations between variational strong convexity and tilt-stable local minimizers.
@@ 1 Introduction
The reason why the classical second-order constructions do not work in the absence of subdifferential continuity, can be simply explained by the fact that for variational convexity we have to work with $f$-attentive localizations of the subgradient mapping whereas the usual limiting coderivative of $\partial f$ does not take $f$-attentive convergence into account. Two possible remedies appear to be manifest: We can either work with generalized derivatives of $f$-attentive localizations of $\partial f$ or we modify the definition of the generalized derivative so that $f$-attentive convergence is incorporated. Although we will show that both approaches are equivalent, the former has the disadvantage that usually we do not know in advance the $f$-attentive neighborhood which we have to work with, whereas the latter results in purely point-based objects.
@@
The paper is organized in the following way. Section 2 contains all the known notions and results of variational analysis and generalized differentiation used in this paper. In Section 3 we introduce the new notion of $f$-attentive generalized derivatives of the subdifferential mapping $\partial f$, provide some elementary calculus and present their relation with generalized derivatives of $f$-attentive localizations of $\partial f$ under prox-regularity. Section 4 provides various neighborhood-based second-order characterizations of variational convexity, whereas in Section 5 we present purely point-based characterizations of variational strong convexity and tilt-stability. Moreover, formulas for the exact bound of the level of variational convexity and tilt-stability are given. The sole assumption we use is prox-regularity, subdifferential continuity is not supposed.
@@ 2 Preliminaries from variational analysis
whenever $x^{\prime}\in{\cal B}_{\epsilon}(\bar{x})$ and $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap({\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}}))$ with $\rho:=f(\bar{x})+\epsilon$, where the set
$\mathrm{gph}_{\rho}\,\partial f:=\{(x,x^{*})\in\mathrm{gph}\,\partial f\,\mid\,f(x)<\rho\}=\mathrm{gph}\,\partial f\cap(L_{f}(\rho)\times\mathbb{R}^{n})$
@@
$f(x^{\prime})\geq f(x)+\langle x^{*},x^{\prime}-x\rangle-\frac{r}{2}\|x^{\prime}-x\|^{2}\quad\mbox{ whenever }x^{\prime}\in U,\ (x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$
@@
If $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is prox-regular at $\bar{x}\in\mathrm{dom}\,f$ for ${\bar{x}^{\ast}}\in\partial f(\bar{x})$, it is well-known that for any twice continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ the function $f+g$ is prox-regular at $\bar{x}$ for the subgradient ${\bar{x}^{\ast}}+\nabla g(\bar{x})\in\partial(f+g)(\bar{x})$, cf. [20, Exercise 13.35]. Subdifferential continuity, if present, is likewise preserved.
@@
$f(x^{\prime})\geq f(x)+\langle x^{*},x^{\prime}-x\rangle+\frac{s}{2}\|x^{\prime}-x\|^{2}\quad\mbox{ whenever }x^{\prime}\in U,\ (x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$ (2.1)
@@
The following proposition is an immediate consequence of a result proven very recently by Khanh, Khoa, Mordukhovich and Phat [9, Theorem 3.3], which is also valid in the infinite dimensional case.
@@
$\langle y^{*}-x^{*},y-x\rangle\geq s\|y-x\|^{2}\quad\mbox{whenever }(x,x^{*}),(y,y^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V).$
@@
$\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\quad\mbox{and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$}.$
@@
The following fundamental equivalence between the properties from Definitions 2.2, 2.5 and 2.6 was very recently shown by Rockafellar [19, Theorem 1].
@@
    Let the function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be variationally $s$-convex at $\bar{x}$ for ${\bar{x}^{\ast}}\in\partial f(\bar{x})$ and let $\widehat{f}$, $U$, $V$ and $\rho$ be as in Definition 2.6. Then for any convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ we also have $\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. Further, for every $\tilde{\rho}\in(f(\bar{x}),\rho]$ there is an open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ with $\mathrm{gph}_{\tilde{\rho}}\,\partial f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. To see this, note that $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$ due to convexity of $\widehat{f}-\frac{s}{2}\|\cdot\|^{2}$. Hence we can find an open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ there holds $f(x)=\widehat{f}(x)<\tilde{\rho}$ and hence
@@
    $=\mathrm{gph}\,\partial f\cap(U\times V)\cap(L_{f}(\rho)\times\mathbb{R}^{n})\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})=\mathrm{gph}_{\tilde{\rho}}\,\partial f\cap(\tilde{U}\times\tilde{V}).$
@@
    $\widehat{f}(x^{\prime}):=\sup\{f(x)+\langle x^{*},x^{\prime}-x\rangle\,\mid\,(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)\}.$
    As the supremum of a collection of affine functions of $x^{\prime}$, $\widehat{f}$ is lsc convex. Further, since $f(x)$ is bounded above by $\rho$ for $(x,x^{*})\in\mathrm{gph}_{\rho}\,\partial f$ and $U\times V$ is bounded, $\widehat{f}$ is everywhere finite and thus strictly continuous. Now a close look at Rockafellar’s proof [16, pp.553–554] of the implication $(c)\Rightarrow(b)$ in [16, Theorem 1] tells us that $\widehat{f}\leq f$ on $U$ and that we can find some open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that $\mathrm{gph}_{\rho}\,\partial f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. Thus our assertion is verified for $s=0$ and for general $s\in\mathbb{R}$ the statement is a consequence of Lemma 4.2 below.
@@
For the subdifferential of a prox-regular and subdifferentially continuous function we have the following result, which can be derived from [2, Proposition 3.6, Lemma 3.4] and gives a certain basis representation of subspaces $L\in{\mathcal{S}}(\partial f)(x,x^{*})$.
@@
Given $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$, a point $\bar{x}\in\mathrm{dom}\,f$ is a tilt-stable local minimizer of $f$ (with modulus $\sigma$) if there exists a number $\gamma>0$ such that the mapping
@@
is single-valued and Lipschitz continuous (with constant $\sigma$) on some neighborhood $V$ of $0\in\mathbb{R}^{n}$ with $M_{\gamma}(0)=\{\bar{x}\}$.
@@
For prox-regular and subdifferentially continuous functions there exist a bunch of characterizations of tilt-stability in the literature, cf. [14, 11, 1]. We present here the following ones.
@@
In this case one has
@@
The equivalence between (i) and (ii) was established in [14]. The equivalence between (i) and (iii) follows from [3, Theorem 7.9] by using the $(P,W)$-basis representation of subspaces $L\in{\mathcal{S}}(\partial f)(\bar{x},0)$. Finally the formulas for ${\rm tilt\;}(f,\bar{x})$ can be found in [3, Theorem 7.9] and [11, Theorem 3.6], respectively. ∎
@@ 3 $f$-attentive derivatives of the subgradient mapping
By elementary calculus rules, cf. [20, Exercise 8.8], we have $\partial h(x)=\partial f(x)+\nabla g(x)$ for all $x$ where $f(x)$ is finite. Now consider a point $(x,x^{*})\in\mathrm{gph}\,\partial f$. Then it is easy to see that for any sequence $(x_{k},x_{k}^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$ there holds $(x_{k},x_{k}^{*})\mathop{\longrightarrow}\limits_{f}^{\mathrm{gph}\,\partial f}(x,x^{*})$ if and only if $(x_{k},x_{k}^{*}+\nabla g(x_{k}))\mathop{\longrightarrow}\limits_{h}^{\mathrm{gph}\,\partial h}(x,x^{*}+\nabla g(x))$. Further, for every sequence $t_{k}\downarrow 0$ there holds $((x_{k},x_{k}^{*})-(x,x^{*}))/t_{k}\to(u,u^{*})$ if and only if
@@
by [15, Corollary 16.3.2]. Since $\lim_{x\to\bar{x}}A(x)$ equals the nonsingular matrix $\bar{A}$, we may also infer that
@@
and, analogously, that $\mathrm{gph}\,D^{*}_{h}(\partial h)(\bar{x},{\bar{x}^{\ast}}+\nabla g(\bar{x}))=\bar{A}\,\mathrm{gph}\,D^{*}_{f}(\partial f)(\bar{x},{\bar{x}^{\ast}})$. Finally, there holds ${\cal O}_{\partial h}^{h}=\{(x,x^{*}+\nabla g(x))\,\mid\,(x,x^{*})\in{\cal O}_{\partial f}^{f}\}$ and ${\mathcal{S}}_{h}(\partial h)(\bar{x},\bar{x}^{*}+\nabla g(\bar{x}))=\{\bar{A}L\,\mid\,L\in{\mathcal{S}}_{f}(\partial f)(\bar{x},{\bar{x}^{\ast}})\}$ follows from [3, Lemma 3.1(iii)]. The formula for the adjoint SC derivative is a consequence of
@@
$\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}=\mathrm{gph}_{f(\bar{x})+\epsilon}\,\partial f\cap({\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})).$
We start our analysis with the following lemma inspired by [13, Proposition 2.3]. Recall that a set $C\subset\mathbb{R}^{n}$ is said to be locally closed at a point $\bar{x}$ (not necessarily in $C$) if $C\cap\bar{\cal B}_{\epsilon}(\bar{x})$ is closed for some $\epsilon>0$.
@@
$\mathrm{gph}_{\rho}\,\partial f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\quad\mbox{and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$.}$
@@
$\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}_{\tilde{\rho}}\,\partial f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\tilde{\cal T}_{\tilde{\epsilon}}^{f}\cap(\tilde{U}\times\tilde{V})$

### v2

@@
Abstract. Variational convexity, together with ist strong counterpart, of extended-real-valued functions has been recently introduced by Rockafellar. In this paper we present second-order characterizations of these properties, i.e., conditions using first-order generalized derivatives of the subgradient mapping. Up to now, such characterizations are only known under the assumptions of prox-regularity and subdifferential continuity and in this paper we discard the latter. To this aim we slightly modify the definitions of the generalized derivatives to be compatible with the $f$-attentive convergence appearing in the definition of subgradients. We formulate our results in terms of both coderivatives and subspace containing derivatives. We also give formulas for the exact bound of variational convexity and study relations between variational strong convexity, tilt-stable local minimizers and strong metric regularity of some truncation of the subgradient mapping.
@@ 1 Introduction
The reason why the classical second-order constructions do not work in the absence of subdifferential continuity, can be simply explained by the fact that for variational convexity we have to work with $f$-attentive localizations of the subgradient mapping whereas the usual limiting coderivative of $\partial f$ does not take $f$-attentive convergence into account. Two possible remedies appear to be manifest: We can either work with generalized derivatives of $f$-attentive localizations of $\partial f$ or we modify the definition of the generalized derivative so that $f$-attentive convergence is incorporated. Although we will show that both approaches are equivalent, the former has the disadvantage that usually we do not know in advance the $f$-attentive neighborhood which we have to work with, whereas the latter results in purely point-based objects. We refer the reader to the very recent paper [11] where a neighborhood-based characterization of variational convexity was given in terms of the regular/limiting coderivative of some $f$-attentive localization of $\partial f$.
@@
The paper is organized in the following way. Section 2 contains all the known notions and results of variational analysis and generalized differentiation used in this paper. In Section 3 we introduce the new notion of $f$-attentive generalized derivatives of the subdifferential mapping $\partial f$, provide some elementary calculus and present their relation with generalized derivatives of $f$-attentive localizations of $\partial f$ under prox-regularity. Section 4 provides various neighborhood-based second-order characterizations of variational convexity, whereas in Section 5 we present purely point-based characterizations of variational strong convexity and tilt-stability and we establish equivalence between these properties and strong meric regularity of a truncation of the subgradient mapping. Moreover, formulas for the exact bound of the level of variational convexity and tilt-stability are given. The sole assumption we use is prox-regularity, subdifferential continuity is not supposed.
@@ 2 Preliminaries from variational analysis
whenever $x^{\prime}\in{\cal B}_{\epsilon}(\bar{x})$ and $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap({\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}}))$ with $\rho:=f(\bar{x})+\epsilon$, where the mapping $\partial_{\rho}f:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ given by
$\mathrm{gph}\,\partial_{\rho}f:=\{(x,x^{*})\in\mathrm{gph}\,\partial f\,\mid\,f(x)<\rho\}=\mathrm{gph}\,\partial f\cap(L_{f}(\rho)\times\mathbb{R}^{n})$
@@
$f(x^{\prime})\geq f(x)+\langle x^{*},x^{\prime}-x\rangle-\frac{r}{2}\|x^{\prime}-x\|^{2}\quad\mbox{ whenever }x^{\prime}\in U,\ (x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$
@@
If $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is prox-regular at $\bar{x}\in\mathrm{dom}\,f$ for ${\bar{x}^{\ast}}\in\partial f(\bar{x})$, it is well-known that for any twice continuously differentiable function $g:\mathbb{R}^{n}\to\mathbb{R}$ the function $f+g$ is prox-regular at $\bar{x}$ for the subgradient ${\bar{x}^{\ast}}+\nabla g(\bar{x})\in\partial(f+g)(\bar{x})$, cf. [22, Exercise 13.35]. Subdifferential continuity, if present, is likewise preserved.
@@
$f(x^{\prime})\geq f(x)+\langle x^{*},x^{\prime}-x\rangle+\frac{s}{2}\|x^{\prime}-x\|^{2}\quad\mbox{ whenever }x^{\prime}\in U,\ (x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$ (2.1)
@@
The following proposition is an immediate consequence of a result proven very recently by Khanh, Khoa, Mordukhovich and Phat [10, Theorem 3.3], which is also valid in the infinite dimensional case.
@@
$\langle y^{*}-x^{*},y-x\rangle\geq s\|y-x\|^{2}\quad\mbox{whenever }(x,x^{*}),(y,y^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V).$
@@
$\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\quad\mbox{and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$}.$
@@
The following fundamental equivalence between the properties from Definitions 2.2, 2.5 and 2.6 was very recently shown by Rockafellar [21, Theorem 1].
@@
    Let the function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be variationally $s$-convex at $\bar{x}$ for ${\bar{x}^{\ast}}\in\partial f(\bar{x})$ and let $\widehat{f}$, $U$, $V$ and $\rho$ be as in Definition 2.6. Then for any convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ we also have $\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. Further, for every $\tilde{\rho}\in(f(\bar{x}),\rho]$ there is an open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ with $\mathrm{gph}\,\partial_{\tilde{\rho}}f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. To see this, note that $\widehat{f}$ is prox-regular and subdifferentially continuous at $\bar{x}$ for ${\bar{x}^{\ast}}$ due to convexity of $\widehat{f}-\frac{s}{2}\|\cdot\|^{2}$. Hence we can find an open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that for every $(x,x^{*})\in\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ there holds $f(x)=\widehat{f}(x)<\tilde{\rho}$ and hence
@@
    $=\mathrm{gph}\,\partial f\cap(U\times V)\cap(L_{f}(\rho)\times\mathbb{R}^{n})\cap(\tilde{U}\times\tilde{V})\cap(L_{f}(\tilde{\rho})\times\mathbb{R}^{n})=\mathrm{gph}\,\partial_{\tilde{\rho}}f\cap(\tilde{U}\times\tilde{V}).$
@@
    $\widehat{f}(x^{\prime}):=\sup\{f(x)+\langle x^{*},x^{\prime}-x\rangle\,\mid\,(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)\}.$
    As the supremum of a collection of affine functions of $x^{\prime}$, $\widehat{f}$ is lsc convex. Further, since $f(x)$ is bounded above by $\rho$ for $(x,x^{*})\in\mathrm{gph}\,\partial_{\rho}f$ and $U\times V$ is bounded, $\widehat{f}$ is everywhere finite and thus strictly continuous. Now a close look at Rockafellar’s proof [18, pp.553–554] of the implication $(c)\Rightarrow(b)$ in [18, Theorem 1] tells us that $\widehat{f}\leq f$ on $U$ and that we can find some open convex neighborhood $\tilde{U}\times\tilde{V}\subset U\times V$ of $(\bar{x},{\bar{x}^{\ast}})$ such that $\mathrm{gph}\,\partial_{\rho}f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})$ and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$. Thus our assertion is verified for $s=0$ and for general $s\in\mathbb{R}$ the statement is a consequence of Lemma 4.2 below.
@@
For the subdifferential of a prox-regular and subdifferentially continuous function we have the following result, which can be derived from [3, Proposition 3.6, Lemma 3.4] and gives a certain basis representation of subspaces $L\in{\mathcal{S}}(\partial f)(x,x^{*})$.
@@
Given $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$, a point $\bar{x}\in\mathrm{dom}\,f$ is a tilt-stable local minimizer of $f$ (with modulus $\kappa$) if there exists a number $\gamma>0$ such that the mapping
@@
is single-valued and Lipschitz continuous (with constant $\kappa$) on some neighborhood $V$ of $0\in\mathbb{R}^{n}$ with $M_{\gamma}(0)=\{\bar{x}\}$.
@@
For prox-regular and subdifferentially continuous functions there exist a bunch of characterizations of tilt-stability in the literature, cf. [16, 1, 13, 2]. We present here the following ones. Recall that a mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is said to be strongly metrically regular with modulus $\kappa$ at $\bar{x}$ for $\bar{y}\in F(\bar{x})$, if $F^{-1}$ has a single-valued localization around $\bar{y}$ for $\bar{x}$ which is Lipschitz continuous with constant $\kappa$.
@@
These statements entail that there is $\delta>0$ such that for all $x^{*}$ in some neighborhood of 00, the mapping $M_{\gamma}$ in Definition 2.11 has $M_{\gamma}(x^{*})$ as the unique $x\in(\partial f)^{-1}(x^{*})$ with $\|x-\bar{x}\|<\delta$ and consequently $\partial f$ is strongly metrically regular at $\bar{x}$ for 00 with modulus equal to the modulus of tilt-stability. Further,
@@
The equivalence between (i) and (ii) together with the identity $M_{\gamma}(x^{*})=(\partial f)^{-1}(x^{*})\cap{\cal B}_{\delta}(\bar{x})$ (without mentioning strong metric regularity of $\partial f$) was established in [16, Theorem 1.3]. The equivalence between (i) and (iii) follows from [4, Theorem 7.9] by using the $(P,W)$-basis representation of subspaces $L\in{\mathcal{S}}(\partial f)(\bar{x},0)$. The formulas for ${\rm tilt\;}(f,\bar{x})$ can be found in [4, Theorem 7.9] and [13, Theorem 3.6], respectively. ∎
@@ 3 $f$-attentive derivatives of the subgradient mapping
By elementary calculus rules, cf. [22, Exercise 8.8], we have $\partial h(x)=\partial f(x)+\nabla g(x)$ for all $x$ where $f(x)$ is finite. Now consider a point $(x,x^{*})\in\mathrm{gph}\,\partial f$. Then it is easy to see that for any sequence $(x_{k},x_{k}^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$ there holds $(x_{k},x_{k}^{*})\mathop{\longrightarrow}\limits_{f}^{\mathrm{gph}\,\partial f}(x,x^{*})$ if and only if $(x_{k},x_{k}^{*}+\nabla g(x_{k}))\mathop{\longrightarrow}\limits_{h}^{\mathrm{gph}\,\partial h}(x,x^{*}+\nabla g(x))$. Further, for every sequence $t_{k}\downarrow 0$ there holds $((x_{k},x_{k}^{*})-(x,x^{*}))/t_{k}\to(u,u^{*})$ if and only if
@@
by [17, Corollary 16.3.2]. Since $\lim_{x\to\bar{x}}A(x)$ equals the nonsingular matrix $\bar{A}$, we may also infer that
@@
and, analogously, that $\mathrm{gph}\,D^{*}_{h}(\partial h)(\bar{x},{\bar{x}^{\ast}}+\nabla g(\bar{x}))=\bar{A}\,\mathrm{gph}\,D^{*}_{f}(\partial f)(\bar{x},{\bar{x}^{\ast}})$. Finally, there holds ${\cal O}_{\partial h}^{h}=\{(x,x^{*}+\nabla g(x))\,\mid\,(x,x^{*})\in{\cal O}_{\partial f}^{f}\}$ and ${\mathcal{S}}_{h}(\partial h)(\bar{x},\bar{x}^{*}+\nabla g(\bar{x}))=\{\bar{A}L\,\mid\,L\in{\mathcal{S}}_{f}(\partial f)(\bar{x},{\bar{x}^{\ast}})\}$ follows from [4, Lemma 3.1(iii)]. The formula for the adjoint SC derivative is a consequence of
@@
$\mathrm{gph}\,\mathcal{T}^{f}_{\epsilon}=\mathrm{gph}\,\partial_{f(\bar{x})+\epsilon}f\cap({\cal B}_{\epsilon}(\bar{x})\times{\cal B}_{\epsilon}({\bar{x}^{\ast}})).$
We start our analysis with the following lemma inspired by [15, Proposition 2.3]. Recall that a set $C\subset\mathbb{R}^{n}$ is said to be locally closed at a point $\bar{x}$ (not necessarily in $C$) if $C\cap\bar{\cal B}_{\epsilon}(\bar{x})$ is closed for some $\epsilon>0$.
@@
$\mathrm{gph}\,\partial_{\rho}f\cap(U\times V)=\mathrm{gph}\,\partial\widehat{f}\cap(U\times V)\quad\mbox{and $f(x)=\widehat{f}(x)$ at the common elements $(x,x^{*})$.}$
@@
$\mathrm{gph}\,\partial\widehat{f}\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\partial_{\tilde{\rho}}f\cap(\tilde{U}\times\tilde{V})=\mathrm{gph}\,\tilde{\cal T}_{\tilde{\epsilon}}^{f}\cap(\tilde{U}\times\tilde{V})$